OpenStudy (anonymous):

what is the midpoint of ab (line on top) with endpoints a(-1,0) and b(5,2)???

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

(2,1)

OpenStudy (anonymous):

x=(-1+5)/2 x=2 y=(2+0)/2 y=1

OpenStudy (anonymous):

\[the\ Midpoint\ is\ (2, 1) \]

OpenStudy (anonymous):

(-1+5)/2 , (0+2)/2 = 2,1

OpenStudy (anonymous):

\[click\ the\ good\ answer\ button\ on\ the\ right\ \]

OpenStudy (anonymous):

all agreed - must be right!